Please explain to me how to answer this question on my graphing calculator.......A company finds that the cost C to produce x items is C=560+12x, while the revenue is R=20.7. Find the break even point. Provide 2 tables to justify your answer.
X C
___ ___
___ ___
___ ___
___ ___
___ ___
X R
___ ___
___ ___
___ ___
___ ___
___ ___
Please help me to understand how to enter this into the calculator....Thank YOu, Tasha

Answer by stanbon(75887)

Please explain to me how to answer this question on my graphing calculator.......A company finds that the cost C to produce x items is C=560+12x, while the revenue is R=20.7. Find the break even point. Provide 2 tables to justify your answer.
Graph the equation y=560+12x
Graph the equation y=20.7
Graph both of these on the same screen.
If you have a TI 8X calculator go to
2nd CALC
Select "Intersect"
Hit the enter button slowly 3 times to see
the point of intersection. Notice the
x value.
Now go to 2nd Table Set
Enter the x value of the intersection point
Then go to 2nd Table.
Notice the values under the columns headed Y1, Y2
Notice the x value where the y values are the same.
Cheers,
Stan H>